### v1

@@ 2.1.4 Structural assumptions
Fourth, the following shifted version of Young’s inequality holds: for all $\delta>0$ there exists a positive constant $C_{\delta}$ such that [15, Lemma 32]
@@ 2.2 The operator $\mathbf{A}$
Let $\Phi$ be an $N$-function that satisfies Assumption 2.1.4. We define the nonlinear vector field $\mathbf{A}$ as follows:
@@
Example 3 (the $p$-Laplacian).
We let $p\in(1,\infty)$ and $\kappa\geq 0$. Recall the function $\Phi_{p,\kappa}$ introduced in (3). Note that $\Phi_{p,\kappa}$ satisfies the conditions in Assumption 2.1.4 and that the characteristics of $\Phi_{p,\kappa}$ do not depend on $\kappa$. In fact, we have
@@
It is shown in [15, Lemma 25] that the $N$-functions $\Psi$ and $\Psi^{*}$ also satisfy Assumption 2.1.4 and that $\Psi^{\prime\prime}(t)\simeq\sqrt{\Phi^{\prime\prime}(t)}$ uniformly in $t>0$. With the function $\Psi$ at hand, we introduce the vector field $\mathbf{V}$ as follows:
@@
We let $p\in(1,\infty)$ and $\kappa\geq 0$. The following functions correspond to those defined in (3):
@@
Proof 2.4.
A proof of the estimates (15)–(18) can be found in [15, Lemma 3] while a proof for (19) can be found in [34, Corollary 64].
2.3 Orlicz spaces
$\varrho(f,\Phi)\coloneqq\int_{\Omega}\Phi(|f(x)|)\mathrm{d}\mu(x).$
The function $\varrho$ is a semimodular on $L^{0}(\Omega)$. Moreover, if $\Phi$ is positive, then $\varrho$ is a modular [16, Lemma 2.3.10]. We define the Orlicz space [40, Sec. 3.1, Definition 5]
equipped with the Luxemburg norm (see [40, Sec. 3.2, (6) and Theorem 3])
(20) $\|f\|_{L^{\Phi}(\mu,\Omega)}\coloneqq\inf\left\{k>0:\varrho\left(\frac{v}{k},\Phi\right)\leq 1\right\}.$
  • $L^{\Phi}(\mu,\Omega)$ is a Banach space [40, Sec. 3.3, Theorem 10].
  • If $\Phi\in\Delta_{2}$, then (see [40, Sec. 3.1, Theorem 2 and Sec. 3.3, Proposition 3])
  • If $\Phi\in\Delta_{2}$, then $L^{\Phi}(\mu,\Omega)$ is separable [40, Sec. 3.5, Thm. 1 and Sec. 3.4, Cor. 5].
  • If $(\Phi,\Phi^{*})$ is a pair of complementary $N$-functions, then we have Hölder’s inequality [40, Sec. 3.3, Proposition 1] (the constant $2$ cannot be omitted):
    $\int_{\Omega}|f(x)g(x)|\mathrm{d}\mu(x)\leq 2\|f\|_{L^{\Phi}(\mu,\Omega)}\|g\|_{L^{\Phi^{*}}(\mu,\Omega)}\qquad\forall f\in L^{\Phi}(\mu,\Omega),\ \forall g\in L^{\Phi^{*}}(\mu,\Omega).$
  • If $(\Phi,\Phi^{*})$ is a pair of complementary $N$-functions with $\Phi,\Phi^{*}\in\Delta_{2}$, then the dual space $[L^{\Phi}(\mu,\Omega)]^{*}$ of $L^{\Phi}(\mu,\Omega)$ is isomorphic to the space $L^{\Phi^{*}}(\mu,\Omega)$ in the following sense [40, Sec. 4.1, Theorem 7]:
    and $\|J_{w}\|_{L^{\Phi}(\mu,\Omega)^{*}}\leq 2\|w\|_{L^{\Phi^{*}}(\mu,\Omega)}$. $L^{\Phi}(\mu,\Omega)$ is reflexive [40, Sec. 4.1, Thm. 10].
We refer the reader to [33, 40, 16, 30, 36] for further facts and properties of Orlicz spaces and their generalizations.
2.4 $A_{p}$ weights
A weight $\omega:\mathbb{R}^{d}\to\mathbb{R}$ is a locally integrable function in $\mathbb{R}^{d}$ which is such that $\omega(x)>0$ for a.e. $x\in\mathbb{R}^{d}$. For a weight $\omega$ and a measurable set $A\subset\mathbb{R}^{d}$, we define $\omega(A)\coloneqq\int_{A}\omega(x)\,\mbox{\rm d}x$. For $p\in[1,\infty)$ we say that a weight $\omega$ belongs to the Muckenhoupt class $A_{p}$ if there is a positive constant $C$ such that for every ball $B\subset\mathbb{R}^{d}$ [43, Definition 1.2.2]
(21) $\left(\fint_{B}\omega(x)\,\mbox{\rm d}x\right)\left(\fint_{B}\omega(x)^{-\tfrac{1}{p-1}}\,\mbox{\rm d}x\right)^{p-1}\leq C,$ $p$ $>1,$
(22) $\left(\fint_{B}\omega(x)\,\mbox{\rm d}x\right)\sup_{x\in B}\frac{1}{\omega(x)}\leq C,$ $p$ $=1.$
The infimum over all such constants $C$ is called the Muckenhoupt characteristic of $\omega$ and is denoted by $[\omega]_{A_{p}}$. We define the class $A_{\infty}$ as $A_{\infty}=\cup_{p\geq 1}A_{p}$. We note that
$\omega\in A_{p}\iff\omega^{\prime}\coloneqq\omega^{-\tfrac{1}{p-1}}\in A_{p^{\prime}},\qquad[\omega]_{A_{p}}=[\omega^{\prime}]_{A_{p^{\prime}}}^{p-1},\qquad p\in(1,\infty);$
see [43, Remark 1.2.4, item 4], where we have denoted the Hölder conjugate of $p$ by $p^{\prime}$. We refer the reader to [43, 19, 32, 28] for more details on Muckenhoupt weights.
Many useful properties follow from the fact that $\omega\in A_{p}$. We mention here one that will be useful for us in the following:
  • Open ended property: If $\omega\in A_{p}$ with $p\in(1,\infty)$, then there is $\delta>0$ such that $\omega\in A_{p-\delta}$; see [43, Corollary 1.2.17] and [19, Corollary 7.6, item (2)].
We refer the reader to [43, 19, 32, 28, 14] for more details on Muckenhoupt weights.
In dealing with the obstacle problem, we restrict ourselves to a class of weights that behave well near the boundary of the domain. The following definition is inspired by [24, Definition 2.5].
Definition 2.5 (class $A_{p}(\Omega)$).
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded Lipschitz domain and let $p\in[1,\infty)$. We say that $\omega\in A_{p}$ belongs to the class $A_{p}(\Omega)$ if there is an open set ${\mathcal{G}}\subset\Omega$, and $\varepsilon,\omega_{l}>0$ such that
$\{x\in\Omega:\operatorname{dist}(x,\partial\Omega)<\varepsilon\}\subset{\mathcal{G}},\qquad\omega_{|\bar{\mathcal{G}}}\in C(\bar{\mathcal{G}}),\qquad\inf_{x\in\bar{\mathcal{G}}}\omega(x)\geq\omega_{l}.$
2.5 Weighted Orlicz spaces
We have finally reached the point where we can describe the functional framework that we will use in our analysis. Namely, that of weighted Orlicz spaces. We let $\omega\in A_{1}$ and define the measure $\mu$ via $\,\mbox{\rm d}\mu(x)=\omega(x)\,\mbox{\rm d}x$. Given an $N$-function $\Phi$, we define the weighted Orlicz space
We endow $L^{\Phi}(\omega,\Omega)$ with the Luxemburg norm defined in (20), and we refer to Section 2.3 for a list of properties of this space.
Let us now assume that $L^{\Phi}(\omega,\Omega)\subset L^{1}_{\mathrm{loc}}(\Omega)$. Then, the elements of $L^{\Phi}(\omega,\Omega)$ are distributions and they have distributional derivatives. We can then define the weighted Orlicz–Sobolev space [36, Definition 3.1.1]
  1. 1.
    $W^{1,\Phi}(\omega,\Omega)$ is a Banach space [30, Theorem 6.1.4(b)].
  2. 2.
    If $\Phi\in\Delta_{2}$, then $W^{1,\Phi}(\omega,\Omega)$ is separable [30, Theorem 6.1.4(c)].
  3. 3.
    If $\Phi,\Phi^{*}\in\Delta_{2}$, then $W^{1,\Phi}(\omega,\Omega)$ is reflexive [30, Theorem 6.1.4(d)].
We define $W_{0}^{1,\Phi}(\omega,\Omega)$ as the closure of $C_{0}^{\infty}(\Omega)\cap W^{1,\Phi}(\omega,\Omega)$ in $W^{1,\Phi}(\omega,\Omega)$ [30, Definition 6.1.8]. It follows that this space has the same properties as $W^{1,\Phi}(\omega,\Omega)$; see [30, Theorem 6.1.9]. In addition, the following modular Poincaré inequality holds.
Proposition 2.6 (modular Poincaré inequality).
Let $\Phi$ be an $N$-function such that $\Phi,\Phi^{*}\in\Delta_{2}$. If $\omega\in A_{1}$, then
Proof 2.7.
    If we set ${\mathcal{F}}\coloneqq\left\{(|w|,|\nabla w|):w\in C_{0}^{\infty}(\Omega)\right\}$, the corresponding inequality is nothing but a weighted Poincaré inequality that holds for every $p_{0}\in(1,\infty)$ and all $w_{0}\in A_{p_{0}}$; see [23, Theorem 1.3].
  • As we do here, [14, Theorem 4.15] assumes that $\Phi$ is an $N$-function with $\Phi,\Phi^{*}\in\Delta_{2}$. We note that in [14] the complementary function of $\Phi$ is denoted by $\bar{\Phi}$.
  • In [14, Theorem 4.15] it is assumed that the weight $\omega$ belongs to $A_{i(\Phi)}$. In our case, we assume that the weight belongs to $A_{1}$ and note that $A_{1}\subset A_{i(\Phi)}$.
The conclusion is thus that (23) holds for all $v\in C_{0}^{\infty}(\Omega)$. Finally, we argue by density. This concludes the proof.
Remark 2.8 (equivalence).
On $W^{1,\Phi}_{0}(\omega,\Omega)$, $\|\nabla w\|_{L^{\Phi}(\omega,\Omega)}$ defines a norm that is equivalent to $\|w\|_{W^{1,\Phi}(\omega,\Omega)}$, provided that $\omega\in A_{1}$ and that $\Phi,\Phi^{*}\in\Delta_{2}$.
To conclude our discussion, we present a suboptimal sufficient condition for the inclusion $L^{\Phi}(\omega,\Omega)\subset L^{1}_{\mathrm{loc}}(\Omega)$.
Let $\Phi$ be an $N$-function. If $\omega\in A_{1}$, then every function in $L^{\Phi}(\omega,\Omega)$ is locally integrable in $\Omega$.
Proof 2.10.
Let $K\subset\Omega$ be compact and let $v\in L^{\Phi}(\omega,\Omega)$. A direct application of the $A_{1}$-condition (22) shows that
$\int_{K}|v(x)|\mathrm{d}x$ $=\int_{K}\omega(x)^{-1}\omega(x)|v(x)|\mathrm{d}x\leq\sup_{x\in K}\frac{1}{\omega(x)}\int_{K}\omega(x)|v(x)|\mathrm{d}x$
$\lesssim\left(\fint_{K}\omega(x)\,\mbox{\rm d}x\right)^{-1}\int_{K}\omega(x)|v(x)|\mathrm{d}x.$
Finally, we control $\int_{K}\omega(x)|v(x)|\mathrm{d}x$ using Young’s inequality (4):
$\int_{K}\omega(x)|v(x)|\mathrm{d}x\leq\int_{K}\omega(x)\left[\Phi(|v(x)|)+\Phi^{*}(1)\right]\mathrm{d}x\\ \lesssim\int_{\Omega}\omega(x)\Phi(|v(x)|)\mathrm{d}x+\Phi^{*}(1)\int_{K}\omega(x)\mathrm{d}x<\infty,$
where we used in the last inequality that a weight is locally integrable in $\mathbb{R}^{d}$.
2.5.1 Weighted Lebesgue and Sobolev spaces
A specific example of the constructions described above are weighted Lebesgue and Sobolev spaces. To introduce them, let $p\in(1,\infty)$ and $\omega\in A_{p}$. The weighted Lebesgue space $L^{p}(\omega,\Omega)$ is obtained by using the $N$-function $\Phi(t)=t^{p}/p$. We immediately note that $L^{p}(\omega,\Omega)\subset L^{1}_{\mathrm{loc}}(\Omega)$ (see [43, bounds (1.2.1)–(1.2.2)]), so we define the weighted Sobolev spaces
$W^{1,p}(\omega,\Omega)\coloneqq\left\{w\in L^{p}(\omega,\Omega):\partial_{i}w\in L^{p}(\omega,\Omega)\,\forall i\in\{1,\ldots,d\}\right\}.$
We define $W_{0}^{1,p}(\omega,\Omega)$ as the closure of $C_{0}^{\infty}(\Omega)$ in $W^{1,p}(\omega,\Omega)$. Due to the fact that $\omega\in A_{p}$ many of the properties of the classical Sobolev spaces extend to the weighted spaces [43, 26, 35]. In particular, we have a weighted Poincaré inequality: if $D\subset\mathbb{R}^{d}$ is open, bounded, and Lipschitz; $p\in(1,\infty)$; and $\omega\in A_{p}$, then [23, Theorem 1.3]
The constant $C_{p}$ depends on $\omega$ only through $[\omega]_{A_{p}}$. The following weighted Poincaré inequality is also useful for our analysis [37, Lemmas 3.1 and 4.2]: There exists a constant $w_{D}$ such that
Finally, we recall a scaled trace inequality. Let $T\subset\mathbb{R}^{d}$ be a simplex and $F\subset T$ a face of $T$. If $p\in[1,\infty)$ and $w\in W^{1,p}(T)$, then its trace $\operatorname{tr}_{\partial T}w\in L^{p}(\partial T)$ satisfies
3 The weighted $\varphi$-Laplace problem
We now present the assumptions that allow us to properly formulate and analyze the problem of minimizing (1). Throughout our discussion, $\Omega\subset\mathbb{R}^{d}$ is a Lipschitz polytope. We let $\omega\in A_{1}$ and let $\varphi$ be an $N$-function that satisfies the conditions in Assumption 2.1.4.
For $f\in L^{\varphi^{*}}(\omega,\Omega)$, the following variational problem is a necessary and sufficient condition for a minimum of $\mathcal{J}$ over $W^{1,\varphi}_{0}(\omega,\Omega)$: Find $u\in W^{1,\varphi}_{0}(\omega,\Omega)$ such that
(27) $\int_{\Omega}\omega(x)\mathbf{A}(\nabla u(x))\cdot\nabla v(x)\mathrm{d}x=\int_{\Omega}\omega(x)f(x)v(x)\mathrm{d}x\quad\forall v\in W_{0}^{1,\varphi}(\omega,\Omega),$
$\left|\int_{\Omega}\omega(x)f(x)v(x)\mathrm{d}x\right|\leq\int_{\Omega}\omega(x)\varphi(|v(x)|)\mathrm{d}x+\int_{\Omega}\omega(x)\varphi^{*}(|f(x)|)\mathrm{d}x<\infty,$
so that the term on the right hand side of (27) is well-defined. A convex minimization argument or the theory of monotone operators guarantee the existence and uniqueness of a weak solution.
3.1 Finite element discretization
Now that we have established the well-posedness of (27), we proceed with its approximation. Since we have assumed that $\Omega$ is a polytope, it can be meshed exactly. We denote by $\mathscr{T}_{h}=\{T\}$ a conforming partition, or mesh, of $\bar{\Omega}$ into closed simplices $T$ of size $h_{T}=\mathrm{diam}(T)$. Here, $h=\max\{h_{T}:T\in\mathscr{T}_{h}\}$. For $T\in\mathscr{T}_{h}$, we define $S_{T}$ as the set of all elements in $\mathscr{T}_{h}$ that share at least one vertex with $T$. We denote by $\mathbb{T}=\{\mathscr{T}_{h}\}_{h>0}$ a collection of conforming meshes that are refinements of an initial mesh $\mathscr{T}_{0}$. We assume that $\mathbb{T}$ satisfies the so-called shape regularity condition [22, Definition 1.107]: there exits a constant $\sigma>1$ such that
@@ 3.1 Finite element discretization
Once again, a convex minimization argument immediately yields the existence and uniqueness of a finite element solution.
@@ 3.2 A best approximation result
We present a best approximation result for the finite element scheme (28). We begin by noting that a direct property of the finite element approximation is Galerkin orthogonality:
@@
Let $u\in W_{0}^{1,\varphi}(\omega,\Omega)$ solve (27), and let $u_{h}\in\mathbb{V}_{h}$ be its finite element approximation defined in (28). Then,
@@
$=\int_{\Omega}\omega(x)(\mathbf{A}(\nabla u(x))-\mathbf{A}(\nabla u_{h}(x)))\cdot\nabla z(x))\mathrm{d}x.$
@@
$\mathfrak{I}$ $\simeq\int_{\Omega}\omega(x)\varphi^{\prime\prime}(|\nabla u(x)|+|\nabla u_{h}(x)|)|\nabla e(x)||\nabla z(x)|\mathrm{d}x$
$\simeq\int_{\Omega}\omega(x)\varphi^{\prime}_{|\nabla u(x)|}(|\nabla e(x)|)|\nabla z(x)|\mathrm{d}x.$
@@
$\mathfrak{I}$ $\leq C\delta\int_{\Omega}\omega(x)\varphi_{|\nabla u(x)|}(|\nabla e(x)|)\mathrm{d}x+C_{\delta}\int_{\Omega}\omega(x)\varphi_{|\nabla u(x)|}(|\nabla z(x)|)\mathrm{d}x$
$\leq C\delta\int_{\Omega}\omega(x)|\mathbf{V}(\nabla u(x))-\mathbf{V}(\nabla u_{h}(x))|^{2}\mathrm{d}x+C_{\delta}\int_{\Omega}\omega(x)\varphi_{|\nabla u(x)|}(|\nabla z(x)|)\mathrm{d}x,$
@@ 3.3 Stability estimates for an interpolation operator
An example of an interpolation operator $\Pi_{h}$ that satisfies these two properties is the so-called Scott–Zhang operator, which was introduced in [42]. The construction of $\Pi_{h}$ is such that it preserves homogeneous boundary conditions, i.e., $v=0$ on $\partial\Omega$ implies $\Pi_{h}v=0$ on $\partial\Omega$ [42, (2.17)]. Moreover, $\Pi_{h}$ is a projection from $W^{1,1}(\Omega)$ to $\mathbb{W}_{h}$ with the property that $W_{0}^{1,1}(\Omega)$ is mapped to $\mathbb{V}_{h}$ [42, Theorem 2.1]. The Scott–Zhang operator also satisfies the stability bound (31). We refer the reader to [42, Theorem 3.1] for a proof of this result and mention that in this case the hidden constant in (31) depends only on the shape-regularity coefficient $\sigma$.
Inspired by the unweighted case treated in [17, Theorem 4.5], we present a stability result in weighted Sobolev–Orlicz spaces $W^{1,\Phi}(\omega,\Omega)$.
@@
Let $\Phi$ be an $N$-function satisfying the conditions in Assumption 2.1.4, let $\omega\in A_{1}$, and let $v\in W^{1,\Phi}(\omega,\Omega)$. Then, for each $T\in\mathscr{T}_{h}$ we have the bound
@@
where the hidden constant depends on $\sigma$, $\Delta_{2}(\Phi)$, and $[\omega]_{A_{1}}$.
Proof 3.4.
@@
The first bound follows from the fact that $(\nabla\Pi_{h}v)_{|K}\in\mathbb{P}_{0}(T)$, while the second estimate follows from the stability of the interpolation operator $\Pi_{h}$ in (31). We note that $C$ depends on the shape-regularity coefficient $\sigma$. We can therefore rely on the monotonicity of $\Phi$ and the $\Delta_{2}$-condition, which holds because $\Phi$ satisfies the conditions in Assumption 2.1.4, to obtain
@@
$\lesssim\fint_{T}\omega(x)\Phi\left(\fint_{S_{T}}|\nabla v(z)|\mathrm{d}z\right)\mathrm{d}x.$
Next, we use Jensen’s inequality and the fact that the weight $\omega$ satisfies the $A_{1}$-condition to derive
$\fint_{T}\omega(x)\Phi(|\nabla\Pi_{h}v(x)|)\mathrm{d}x\lesssim\fint_{T}\omega(x)\fint_{S_{T}}\Phi(|\nabla v(z)|)\mathrm{d}z\mathrm{d}x\\ \leq C\fint_{T}\omega(x)\mathrm{d}x\left(\sup_{z\in S_{T}}\frac{1}{\omega(z)}\right)\fint_{S_{T}}\omega(z)\Phi(|\nabla v(z)|)\mathrm{d}z.$
Since $T\subset S_{T}$ with $|T|\simeq|S_{T}|$, an application of the $A_{1}$-condition (22) allows us to conclude the desired stability estimate.
Next, we present the result of Theorem 32 in the context of a family of shifted $N$-functions.
@@
Let $\omega\in A_{1}$, and let $\{\Phi_{a}\}_{a\geq 0}$ be the family of shifted $N$-functions defined in (7) associated with the $N$-function $\Phi$ that satisfies the conditions in Assumption 2.1.4. Then, we have the following stability bound uniformly in $a\geq 0$ and $T\in\mathscr{T}_{h}$:
@@
where the hidden constant depends on $\sigma$, $\Delta_{2}(\Phi)$, and $[\omega]_{A_{1}}$.
Proof 3.6.
The desired bound results from the application of the stability bound of Theorem 32 to the family of shifted $N$-functions in conjunction with the fact that such a family satisfies the $\Delta_{2}$-condition uniformly in $a\geq 0$ with a constant that depends only on $\Delta_{2}(\Phi)$; see [15, Lemma 23] and [17, Lemma 6.1].

### v2

@@ 3.1 Some further properties
The following shifted version of Young’s inequality holds: for all $\delta>0$ there exists a positive constant $C_{\delta}$ such that [15, Lemma 32]
@@ 3.2 The operator $\mathbf{A}$
Let $\Phi$ be an $N$-function that satisfies Assumption 3. We recall the definition of the vector field $\mathbf{A}$ in (2), namely,
@@
Example 3.1 (the $p$-Laplacian).
Let $p\in(1,\infty)$, and let $\kappa\geq 0$. Recall the function $\Phi_{p,\kappa}$ introduced in (6). Note that $\Phi_{p,\kappa}$ satisfies the conditions in Assumption 3 and that the characteristics of $\Phi_{p,\kappa}$ do not depend on $\kappa$. In fact, we have
@@
In [15, Lemma 25] it is shown that the $N$-functions $\Psi$ and $\Psi^{*}$ also satisfy Assumption 3 and that $\Psi^{\prime\prime}(t)\simeq\sqrt{\Phi^{\prime\prime}(t)}$ uniformly in $t>0$. With the function $\Psi$ at hand, we introduce the vector field $\mathbf{V}$ as follows:
@@
Let $p\in(1,\infty)$, and let $\kappa\geq 0$. The following functions correspond to those defined in (6):
@@
Proof 3.5.
A proof of the estimates (30)–(33) can be found in [15, Lemma 3] while a proof for (34) can be found in [37, Corollary 64].
4 Finite element discretization of the $\Phi$-Laplace problem
Now that we have established the well-posedness of (23), we proceed with its approximation. Since we have assumed that $\Omega$ is a polytope, it can be meshed exactly. Denote by $\mathscr{T}_{h}=\{T\}$ a conforming partition, or mesh, of $\bar{\Omega}$ into closed simplices $T$ of size $h_{T}=\mathrm{diam}(T)$. Here, $h=\max\{h_{T}:T\in\mathscr{T}_{h}\}$. For $T\in\mathscr{T}_{h}$, we define $S_{T}$ as the set of all elements in $\mathscr{T}_{h}$ that share at least one vertex with $T$. We denote by $\mathbb{T}=\{\mathscr{T}_{h}\}_{h>0}$ a collection of conforming meshes that are refinements of an initial mesh $\mathscr{T}_{0}$. We assume that $\mathbb{T}$ satisfies the so-called shape regularity condition [22, Definition 1.107]: there exits a constant $\sigma>1$ such that
@@ 4 Finite element discretization of the $\Phi$-Laplace problem
A convex minimization argument yields the existence and uniqueness of $u_{h}$.
@@ 4.1 A best approximation result
We present a best approximation result for (35). To derive it, we first establish the following Galerkin orthogonality property:
@@
Let $u\in W_{0}^{1,\Phi}(\omega,\Omega)$ solve (23), and let $u_{h}\in\mathbb{V}_{h}$ be its finite element approximation defined as the solution to (35). Then,
@@
$=\int_{\Omega}\omega(x)(\mathbf{A}(\nabla u(x))-\mathbf{A}(\nabla u_{h}(x)))\cdot\nabla z(x)\mathrm{d}x.$
@@
$\mathfrak{I}$ $\simeq\int_{\Omega}\omega(x)\Phi^{\prime\prime}(|\nabla u(x)|+|\nabla u_{h}(x)|)|\nabla e(x)||\nabla z(x)|\mathrm{d}x$
$\simeq\int_{\Omega}\omega(x)\Phi^{\prime}_{|\nabla u(x)|}(|\nabla e(x)|)|\nabla z(x)|\mathrm{d}x.$
@@
$\mathfrak{I}$ $\leq C\delta\int_{\Omega}\omega(x)\Phi_{|\nabla u(x)|}(|\nabla e(x)|)\mathrm{d}x+C_{\delta}\int_{\Omega}\omega(x)\Phi_{|\nabla u(x)|}(|\nabla z(x)|)\mathrm{d}x$
$\leq C\delta\int_{\Omega}\omega(x)|\mathbf{V}(\nabla u(x))-\mathbf{V}(\nabla u_{h}(x))|^{2}\mathrm{d}x+C_{\delta}\int_{\Omega}\omega(x)\Phi_{|\nabla u(x)|}(|\nabla z(x)|)\mathrm{d}x,$
@@ 4.2 Stability estimates for an interpolation operator
An example of an interpolation operator $\Pi_{h}$ that satisfies these two properties is the so-called Scott–Zhang operator, which was introduced in [45]. The construction of $\Pi_{h}$ is such that it preserves homogeneous boundary conditions, i.e., $v=0$ on $\partial\Omega$ implies $\Pi_{h}v=0$ on $\partial\Omega$ [45, (2.17)]. Moreover, $\Pi_{h}$ is a projection from $W^{1,1}(\Omega)$ to $\mathbb{W}_{h}$ with the property that $W_{0}^{1,1}(\Omega)$ is mapped to $\mathbb{V}_{h}$ [45, Theorem 2.1]. The Scott–Zhang operator also satisfies the stability bound (38). We refer the reader to [45, Theorem 3.1] for a proof of this result and mention that in this case the hidden constant in (38) depends only on the shape-regularity coefficient $\sigma$.
Inspired by the unweighted case discussed in [17, Theorem 4.5], we present a stability result in weighted Orlicz–Sobolev spaces $W^{1,\Phi}(\omega,\Omega)$.
@@
Let $\Phi$ be an $N$-function that satisfies Assumption 3, let $\omega\in A_{\Phi}$, and let $v\in W^{1,\Phi}(\omega,\Omega)$. Then, for $T\in\mathscr{T}_{h}$, we have
@@
where the hidden constant depends on $\sigma$, $\omega]_{A_{i(\Phi)}}$, and the characteristics of $\Phi$.
Proof 4.4.
@@
The first bound follows from $(\nabla\Pi_{h}v)_{|T}\in\mathbb{P}_{0}(T)$, while the second follows from the stability of the interpolation operator $\Pi_{h}$ in (38). Note that $C$ depends only on the shape-regularity coefficient $\sigma$. We can therefore rely on the monotonicity of $\Phi$ and the $\Delta_{2}$-condition, which holds because $\Phi$ satisfies Assumption 3, to obtain
@@
$\lesssim\fint_{T}\omega(x)\Phi\left(\fint_{S_{T}}|\nabla v(z)|\mathrm{d}z\right)\mathrm{d}x$
$=\fint_{T}\omega(x)\Phi\left(\fint_{S_{T}}|\nabla v(z)\chi_{S_{T}}(z)|\,\mbox{\rm d}z\right)\,\mbox{\rm d}x$
$\lesssim\fint_{T}\omega(x)\Phi\left(M[\nabla v\chi_{S_{T}}](x)\right)\,\mbox{\rm d}x.$
In the last step, we used shape regularity to replace the average on the patch $S_{T}$ by the average over a ball $B$, which is such that $x\in B$, $S_{T}\subset B$, and $|S_{T}|\approx|B|$. The average over the ball $B$ is therefore bounded by the maximal function. Since $\omega\in A_{\Phi}$, we then invoke Theorem 2.5 to obtain
$\fint_{T}\omega(x)\Phi(|\nabla\Pi_{h}v(x)|)\mathrm{d}x$ $\lesssim\frac{1}{|T|}\int_{\mathbb{R}^{d}}\omega(x)\Phi\left(M[\nabla v\chi_{S_{T}}](x)\right)\,\mbox{\rm d}x$
$\lesssim\frac{1}{|T|}\int_{\mathbb{R}^{d}}\omega(x)\Phi\left(|\nabla v(x)\chi_{S_{T}}(x)|\right)\,\mbox{\rm d}x$
$\lesssim\frac{1}{|S_{T}|}\int_{S_{T}}\omega(x)\Phi\left(|\nabla v(x)|\right)\,\mbox{\rm d}x=\fint_{S_{T}}\omega(x)\Phi\left(|\nabla v(x)|\right)\,\mbox{\rm d}x.$
Note that in the course of our derivations all constants depend only on $\sigma$, $\omega]_{A_{i(\Phi)}}$, and the characteristics of $\Phi$. This concludes the proof.
We present the result of Theorem 39 in the context of a family of shifted functions.
@@
Let $\Phi$ be an $N$-function satisfying Assumption 3, and let $\{\Phi_{a}\}_{a\geq 0}$ be the family of shifted $N$-functions defined in (10) associated with $\Phi$. If $\omega\in A_{\Phi}$, then we have the following bound uniformly in $a\geq 0$ and $T\in\mathscr{T}_{h}$:
@@
where the hidden constant depends on $\sigma$, $\omega]_{A_{i(\Phi)}}$, and the characteristics of $\Phi$.
Proof 4.6.
The desired bound results from the application of the stability bound of Theorem 39 to the family of shifted $N$-functions $\{\Phi_{a}\}_{a\geq 0}$ in conjunction with the fact that such a family satisfies the $\Delta_{2}$-condition uniformly in $a\geq 0$ with a constant that depends only on $\Delta_{2}(\Phi)$; see [15, Lemma 23] and [17, Lemma 6.1].